# A unified approach to hypergeometric class functions

### v1

@@ 3.4 Basic pair of factorizations
$\begin{array}[]{rl}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}&=\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}\partial_{z}+n(n+1)\frac{\sigma^{\prime\prime}}{2}+(n+\frac{1}{2})\kappa_{0}^{\prime}+\omega_{0}\\[8.53581pt] &=\partial_{z}\Big{(}\sigma(z)\partial_{z}+\kappa_{n}(z)\Big{)}+n(n-1)\frac{\sigma^{\prime\prime}}{2}+(n-\frac{1}{2})\kappa_{0}^{\prime}+\omega_{0}.\end{array}$
@@
$\begin{array}[]{rl}\partial_{z}\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}&=\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\big{)}\partial_{z},\\[8.53581pt] \Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}&=\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}.\end{array}$
@@
$\begin{array}[]{rll}&\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}F=0&\Rightarrow\qquad\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}\partial_{z}F=0,\\[8.53581pt] &\Big{(}{\cal C}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}F=0&\Rightarrow\qquad\Big{(}{\cal C}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}F=0.\end{array}$
@@
Suppose that for $n\in{\mathbb{C}}$, ${\cal A}_{+}^{n+\frac{1}{2}}$, ${\cal A}_{-}^{n-\frac{1}{2}}$, ${\cal C}^{n}$ are operators and $\eta_{n\pm\frac{1}{2}}$ are numbers satisfying
$\begin{array}[]{rl}{\cal C}^{n}&={\cal A}_{+}^{n+\frac{1}{2}}{\cal A}_{-}^{n-\frac{1}{2}}+\eta_{n-\frac{1}{2}},\\[8.53581pt] &={\cal A}_{-}^{n-\frac{1}{2}}{\cal A}_{+}^{n+\frac{1}{2}}+\eta_{n+\frac{1}{2}}.\end{array}$
@@
$\begin{array}[]{l}{\cal A}_{-}^{n-\frac{1}{2}}{\cal C}^{n}={\cal C}^{n-1}{\cal A}_{-}^{n-\frac{1}{2}},\\[8.53581pt] {\cal A}_{+}^{n+\frac{1}{2}}{\cal C}^{n}={\cal C}^{n+1}{\cal A}_{+}^{n+\frac{1}{2}}\end{array}$
@@
$\begin{array}[]{l}{\cal A}_{+}^{n+\frac{1}{2}}=\partial_{z},\ \ \ \ {\cal A}_{-}^{n-\frac{1}{2}}=\sigma(z)\partial_{z}+\kappa_{n}(z),\ \ \ \ \eta_{n+\frac{1}{2}}=n(n+1)\frac{\sigma^{\prime\prime}}{2}+(n+\frac{1}{2})\kappa^{\prime}+\omega,\\[8.53581pt] {\cal C}^{n}:={\mathcal{C}}(\sigma,\kappa_{n})+\omega_{n}.\end{array}$ (3.13)
@@
Note that ${\mathcal{C}}(\sigma,\kappa_{n})+\omega_{n}$ from this section essentially coincides with ${\mathcal{C}}+\omega$ restricted to ${\cal Z}^{n}$ from Section 2.
@@ 3.5 Ladders of solutions
Let $\kappa_{n},\omega_{n}$ be as in (3.11) and (3.12). Suppose we have a solution
$\Big{(}{\mathcal{C}}(\sigma,\kappa_{0})+\omega_{0}\Big{)}f=0.$ (3.14)
@@
${\mathcal{C}}(\sigma,\kappa_{n})+\omega_{n}\quad\text{annihilates}\quad$ $\partial_{z}^{n}f,$ (3.15)
${\mathcal{C}}(\sigma,\kappa_{-n})+\omega_{-n}\quad\text{annihilates}\quad$ $\sigma^{n}\rho_{0}^{-1}\partial_{z}^{n}\rho_{0}f.$ (3.16)
Note that (3.15) is related to (3.16) by the basic symmetry (3.10).
@@
$\sigma(z)\partial_{z}+\kappa_{-j}(z)$ $=\sigma(z)^{j+1}\rho(z)^{-1}\partial_{z}^{j}\sigma(z)^{-j}\rho(z),$ (3.17)
@@
${\mathcal{C}}(\sigma,\kappa_{0})-\frac{\kappa_{0}^{\prime}}{2}$ $=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa_{0}(z)\big{)}\partial_{z},$ (3.19)
${\mathcal{C}}(\sigma,\kappa_{0})+\frac{\kappa_{0}^{\prime}}{2}$ $=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa_{0}(z)\big{)}\partial_{z}+\kappa_{0}^{\prime}.$ (3.20)
They have elementary solutions
${\mathcal{C}}(\sigma,\kappa_{0})-\frac{\kappa_{0}^{\prime}}{2}\qquad\text{ annihilates }$ $1,$ (3.21)
${\mathcal{C}}(\sigma,\kappa_{0})+\frac{\kappa_{0}^{\prime}}{2}\qquad\text{ annihilates }$ $\rho_{0}^{-1}.$ (3.22)
@@
Alternatively, we can derive (3.22) from (3.21) by the symmetry (3.10). The special solutions (3.21) and (3.22) lead to a pair of one-sided ladders linked by the symmetry (3.10): for any $n\in{\mathbb{N}}_{0}$
${\mathcal{C}}(\sigma,-n\sigma^{\prime}+\kappa_{0})-(n+\tfrac{1}{2})\kappa_{0}^{\prime}+n^{2}\frac{\sigma^{\prime\prime}}{2}\quad\text{annihilates}\quad$ $\sigma^{n}\rho_{0}^{-1}\partial_{z}^{n}\rho_{0}$ (3.24)
${\mathcal{C}}(\sigma,n\sigma^{\prime}+\kappa_{0})+(n+\tfrac{1}{2})\kappa_{0}^{\prime}+n^{2}\frac{\sigma^{\prime\prime}}{2}\quad\text{annihilates}\quad$ $\partial_{z}^{n}\rho_{0}^{-1}.$ (3.25)
(3.24) consists of polynomials, which we will analyze in more detail in Section 7.
@@ 3.6 Chebyshev ladder
${\mathcal{C}}\Big{(}\sigma,-\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $=\sqrt{\sigma(z)}\partial_{z}\sqrt{\sigma(z)}\partial_{z}+\omega$ (3.26)
@@
${\mathcal{C}}\Big{(}\sigma,\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $=\partial_{z}\sqrt{\sigma(z)}\partial_{z}\sqrt{\sigma(z)}+\omega$ (3.28)
@@
${\mathcal{C}}\Big{(}\sigma,-\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $\qquad\text{annihilates}\qquad\Big{(}A\sin\big{(}\omega y(z)\big{)}+B\cos\big{(}\omega y(z)\big{)}\Big{)},$ (3.31)
${\mathcal{C}}\Big{(}\sigma,\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $\qquad\text{annihilates}\qquad\frac{1}{\sqrt{\sigma(z)}}\Big{(}A\sin\big{(}\omega y(z)\big{)}+B\cos\big{(}\omega y(z)\big{)}\Big{)}.$ (3.32)
We can embed the parameters of (3.26) and (3.28) into a single ladder as follows. We set $\kappa_{0}=0$ and $\omega_{0}=\omega-\frac{\sigma^{\prime\prime}}{8}$, so that
$\kappa_{m}=m\sigma^{\prime},\quad\omega_{m}=\omega+\Big{(}m^{2}-\frac{1}{4}\Big{)}\frac{\sigma^{\prime\prime}}{2}.$ (3.33)
@@
By using the recurrence relations (3.15) and (3.16) we obtain elementary solutions for ${\mathcal{C}}(\sigma,\kappa_{m})+\omega_{m}$ with $m\in{\mathbb{Z}}+\frac{1}{2}$.
@@ 4 Singular point
We will discuss solutions of hypergeometric class equations given by power series around 00. We will also describe some symmetries, which exist if (4.1) holds.
@@ 4.1 Solutions around a singular point
$\frac{1}{\sigma(z)}\big{(}{\mathcal{C}}+\omega\big{)}f(z)=\Big{(}\partial_{z}^{2}+\frac{\tau(z)}{\sigma(z)}\partial_{z}+\frac{\omega}{\sigma(z)}\Big{)}f(x)=0$ (4.2)
@@
$\Big{(}{\cal C}(\sigma,\kappa)+\omega\Big{)}F(z)=0,\ \ \ F(0)=1.$ (4.3)
@@
If $\sigma^{\prime}(0)=0$ and $\sigma^{\prime\prime}\neq 0$, then $\sigma(z)$ has a double root at zero and then the series (4.4) is divergent. However, it can be shown to be asymptotic to one of the solutions of (3.6) (see Section 6.3). In all other cases the series (4.4) is convergent. The function $F(\sigma,\kappa,\omega;z)$ given by (4.4) will be called the unified hypergeometric function. It depends meromorphically on $\sigma^{\prime}(0),\sigma^{\prime\prime},\kappa,\omega$.
@@ 4.2 Olver’s normalization
Assume now $\sigma(0)\neq 0$. We set $m:=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. Without sacrificing the generality we can suppose additionally that $\sigma^{\prime}(0)=1$. Thus
@@
00 is a regular singular point and 00, $-m$ are its indices [WW].
@@
Note that ${\mathbf{F}}$ is holomorphic in $\sigma^{\prime\prime},\kappa,\omega$. Dividing by $\Gamma(1+m)$, as in (4.8), is sometimes called Olver’s normalization. Here is the basic pair of recurrence relations for the Olver normalized unified hypergeometric function:

### v2

@@ 3.4 Basic pair of factorizations
Note that ${\cal H}(\sigma,\kappa_{n})+\omega_{n}$ coincides with ${\mathcal{C}}^{n}+\omega_{0}$, where ${\mathcal{C}}^{n}$ is the Casimir for $\sigma,\kappa_{0}$ restricted to $N=n$, see (2.41).
@@
$\begin{array}[]{rl}{{\cal H}}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}&=\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}\partial_{z}+n(n+1)\frac{\sigma^{\prime\prime}}{2}+(n+\frac{1}{2})\kappa_{0}^{\prime}+\omega_{0}\\[8.53581pt] &=\partial_{z}\Big{(}\sigma(z)\partial_{z}+\kappa_{n}(z)\Big{)}+n(n-1)\frac{\sigma^{\prime\prime}}{2}+(n-\frac{1}{2})\kappa_{0}^{\prime}+\omega_{0}.\end{array}$
@@
$\begin{array}[]{rl}\partial_{z}\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}&=\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\big{)}\partial_{z},\\[8.53581pt] \Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}&=\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}.\end{array}$
@@
$\begin{array}[]{rll}&\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}F=0&\Rightarrow\qquad\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}\partial_{z}F=0,\\[8.53581pt] &\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n+1}\big{)}+\omega_{n+1}\Big{)}F=0&\Rightarrow\qquad\Big{(}{{\cal H}}\big{(}\sigma,\kappa_{n}\big{)}+\omega_{n}\Big{)}\Big{(}\sigma(z)\partial_{z}+\kappa_{n+1}(z)\Big{)}F=0.\end{array}$
@@
Suppose that for $n\in{\mathbb{C}}$, ${\cal A}_{+}^{n+\frac{1}{2}}$, ${\cal A}_{-}^{n-\frac{1}{2}}$, ${{\cal H}}^{n}$ are operators and $\eta_{n\pm\frac{1}{2}}$ are numbers satisfying
$\begin{array}[]{rl}{{\cal H}}^{n}&={\cal A}_{+}^{n+\frac{1}{2}}{\cal A}_{-}^{n-\frac{1}{2}}+\eta_{n-\frac{1}{2}},\\[8.53581pt] &={\cal A}_{-}^{n-\frac{1}{2}}{\cal A}_{+}^{n+\frac{1}{2}}+\eta_{n+\frac{1}{2}}.\end{array}$
@@
$\begin{array}[]{l}{\cal A}_{-}^{n-\frac{1}{2}}{{\cal H}}^{n}={{\cal H}}^{n-1}{\cal A}_{-}^{n-\frac{1}{2}},\\[8.53581pt] {\cal A}_{+}^{n+\frac{1}{2}}{{\cal H}}^{n}={{\cal H}}^{n+1}{\cal A}_{+}^{n+\frac{1}{2}}.\end{array}$
@@
$\begin{array}[]{l}{\cal A}_{+}^{n+\frac{1}{2}}=\partial_{z},\ \ \ \ {\cal A}_{-}^{n-\frac{1}{2}}=\sigma(z)\partial_{z}+\kappa_{n}(z),\ \ \ \ \eta_{n+\frac{1}{2}}=n(n+1)\frac{\sigma^{\prime\prime}}{2}+(n+\frac{1}{2})\kappa^{\prime}+\omega,\\[8.53581pt] {{\cal H}}^{n}:={{\cal H}}(\sigma,\kappa_{n})+\omega_{n}.\end{array}$ (3.15)
@@ 3.5 Ladders of solutions
Let $\kappa_{n},\omega_{n}$ be as in (3.13) and (3.14). Let $\rho_{0}$ solve (3.8) for $\sigma_{0},\kappa_{0}$, that is,
$\big{(}\sigma_{0}(z)\partial_{z}-\kappa_{0}(z)\big{)}\rho_{0}(z)=0.$ (3.16)
Suppose we have a solution
$\Big{(}{{\cal H}}(\sigma,\kappa_{0})+\omega_{0}\Big{)}f=0.$ (3.17)
@@
${{\cal H}}(\sigma,\kappa_{n})+\omega_{n}\quad\text{annihilates}\quad$ $\partial_{z}^{n}f,$ (3.18)
${{\cal H}}(\sigma,\kappa_{-n})+\omega_{-n}\quad\text{annihilates}\quad$ $\sigma^{n}\rho_{0}^{-1}\partial_{z}^{n}\rho_{0}f.$ (3.19)
@@
$\sigma(z)\partial_{z}+\kappa_{-j}(z)$ $=\sigma(z)^{j+1}\rho_{0}(z)^{-1}\partial_{z}^{j}\sigma(z)^{-j}\rho_{0}(z),$ (3.20)
@@
${{\cal H}}(\sigma,\kappa_{0})-\frac{\kappa_{0}^{\prime}}{2}$ $=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa_{0}(z)\big{)}\partial_{z},$ (3.22)
${{\cal H}}(\sigma,\kappa_{0})+\frac{\kappa_{0}^{\prime}}{2}$ $=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa_{0}(z)\big{)}\partial_{z}+\kappa_{0}^{\prime}.$ (3.23)
They have elementary solutions:
${{\cal H}}(\sigma,\kappa_{0})-\frac{\kappa_{0}^{\prime}}{2}\qquad\text{ annihilates }$ $1,$ (3.24)
${{\cal H}}(\sigma,\kappa_{0})+\frac{\kappa_{0}^{\prime}}{2}\qquad\text{ annihilates }$ $\rho_{0}^{-1}.$ (3.25)
@@
Alternatively, we can derive (3.25) from (3.24) by the basic symmetry (3.10). Indeed, (3.24) and (3.10) imply that ${\cal H}(\sigma,-\kappa_{0})-\frac{\kappa_{0}^{\prime}}{2}$ annihilates $\rho_{0}$. Then we switch the sign of $\kappa_{0}$, which corresponds to replacing $\rho_{0}$ with $\rho_{0}^{-1}$.
The special solutions (3.24) and (3.25) lead to a pair of one-sided ladders: for any $n\in{\mathbb{N}}_{0}$
${{\cal H}}(\sigma,-n\sigma^{\prime}+\kappa_{0})-(n+\tfrac{1}{2})\kappa_{0}^{\prime}+n^{2}\frac{\sigma^{\prime\prime}}{2}\quad\text{annihilates}\quad$ $\sigma^{n}\rho_{0}^{-1}\partial_{z}^{n}\rho_{0}$ (3.27)
${{\cal H}}(\sigma,n\sigma^{\prime}+\kappa_{0})+(n+\tfrac{1}{2})\kappa_{0}^{\prime}+n^{2}\frac{\sigma^{\prime\prime}}{2}\quad\text{annihilates}\quad$ $\partial_{z}^{n}\rho_{0}^{-1}.$ (3.28)
(3.27) consists of polynomials, which we will analyze in more detail in Section 7. The ladder (3.28) consists of functions of the form $\rho_{0}^{-1}\sigma^{-n}P_{n}$, where $P_{n}:=\sigma^{n}\rho_{0}\partial_{z}^{n}\rho_{0}^{-1}$ are polynomials.
@@ 3.6 Chebyshev ladder
${{\cal H}}\Big{(}\sigma,-\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $=\sqrt{\sigma(z)}\partial_{z}\sqrt{\sigma(z)}\partial_{z}+\omega$ (3.29)
@@
${{\cal H}}\Big{(}\sigma,\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $=\partial_{z}\sqrt{\sigma(z)}\partial_{z}\sqrt{\sigma(z)}+\omega$ (3.31)
@@
${{\cal H}}\Big{(}\sigma,-\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $\qquad\text{annihilates}\qquad\Big{(}A\sin\big{(}\omega y(z)\big{)}+B\cos\big{(}\omega y(z)\big{)}\Big{)},$ (3.34)
${{\cal H}}\Big{(}\sigma,\frac{\sigma^{\prime}}{2}\Big{)}+\omega+\frac{\sigma^{\prime\prime}}{4}$ $\qquad\text{annihilates}\qquad\frac{1}{\sqrt{\sigma(z)}}\Big{(}A\sin\big{(}\omega y(z)\big{)}+B\cos\big{(}\omega y(z)\big{)}\Big{)}.$ (3.35)
We can embed the parameters of (3.29) and (3.31) into a single ladder as follows. We set $\kappa_{0}=0$ and $\omega_{0}=\omega{+}\frac{\sigma^{\prime\prime}}{8}$, so that
$\kappa_{m}=m\sigma^{\prime},\quad\omega_{m}=\omega+\Big{(}m^{2}{+}\frac{1}{4}\Big{)}\frac{\sigma^{\prime\prime}}{2}.$ (3.36)
@@
By using the recurrence relations (3.18) and (3.19) we obtain elementary solutions for
${\cal H}(\sigma,\kappa_{m})+\omega_{m}=\sigma(z)\partial_{z}^{2}+(m+1)\sigma^{\prime}(z)\partial_{z}+\Big{(}m+\frac{1}{2}\Big{)}^{2}\frac{\sigma^{\prime\prime}}{2}+\omega$ (3.39)
with $m\in{\mathbb{Z}}+\frac{1}{2}$.
Remark 3.4
Chebyshev ladders are very special, since $\kappa$ is determined by $\sigma$. The only two nontrivial hypergeometric types where they appear are ${}_{2}F_{1}$ and $F_{1}$. However, they are quite important. One can note that Bessel functions of half-integer parameters are elementary functions because they correspond to the Chebyshev ladder of the $F_{1}$ type.
@@ 4 Singular point
We will discuss solutions of hypergeometric class equations given by power series around 00. We will also describe some symmetries which exist if (4.1) holds.
@@ 4.1 Solutions around a singular point
$\frac{1}{\sigma(z)}\big{(}{\sigma(z)\partial_{z}^{2}+\tau(z)\partial_{z}}+\omega\big{)}f(z)=\Big{(}\partial_{z}^{2}+\frac{\tau(z)}{\sigma(z)}\partial_{z}+\frac{\omega}{\sigma(z)}\Big{)}f(x)=0$ (4.2)
@@
$\Big{(}{{\cal H}}(\sigma,\kappa)+\omega\Big{)}F(z)=0,\ \ \ F(0)=1.$ (4.3)
@@
If $\sigma^{\prime}(0)=0$ and $\sigma^{\prime\prime}\neq 0$, then $\sigma(z)$ has a double root at zero and then the series (4.4) either terminates and defines a polynomial, or is divergent. However, it is always asymptotic to one of the solutions of (3.6) defined on \${\mathbb{C}}\backslash[0,\infty[$ (see Section 6.3 and Appendix A). In all other cases the series (4.4) has a nonzero radius of convergence. The function $F(\sigma,\kappa,\omega;z)$ given by (4.4) will be called the unified hypergeometric function. It depends meromorphically on $\sigma^{\prime}(0),\sigma^{\prime\prime},\kappa,\omega$.
@@ 4.2 Olver’s normalization
Assume now $\sigma^{\prime}(0)\neq 0$. We set $m:=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. Without sacrificing the generality we can suppose additionally that $\sigma^{\prime}(0)=1$. Thus
@@
00 is a regular singular point and 00, $-m$ are its indices, see [WW] or Appendix B.
@@
Note that ${\mathbf{F}}$ is holomorphic in $\sigma^{\prime\prime},\kappa,\omega$. Dividing by $\Gamma(1+m)$, as in (4.8), is sometimes called Olver’s normalization, since it was made popular by Frank Olver’s textbook [Ol2]. Here is the basic pair of recurrence relations for the Olver normalized unified hypergeometric function: